Someone witnesses a crime involving a taxi. The witness tells the police that the taxi was blue. The police know that 89% of witnesses are correct and also that 17% of all of the taxis that could have been involved are blue while the rest are green. What is the probability that the taxi was in fact blue? Write your answer out to two decimal places so that 1% is .01, 99% is .99, etc.
Answer #1

1) WRITE YOUR ANSWER OUT TO TWO DECIMAL PLACES SO THAT 1% IS .01 99% IS ,99?

A)Its classic case of finding the probability of the car.let us assume that there are two cars with unknown value that is x and they are GREEN and BLUE.we need to find out the probability of blue being taken care of out of the x.so according to the police investigation they came to know that the color of the car is BLUE and out of 100 they think that 17 are blue so this can be x-17 blue cars.and they out of 100 the second probability will be green so it will be green that is x it is 100 so from the equation it states that 100-17=83 so green cars are 83 and they are 17 blue cars.now lets assume that by the case of the police that 89% of the witness say the truth so the probability will be 0.89 out of 1 so can be given a nod that the witness given the correct answer so lets move on that so out of 17 blue cars that out of 100 they will be 100/17=5.88,now out of this 83/17=4.88 so the answer will be 5.88/4.88=1.20.